OpenStudy (anonymous):

5i/(2+3i)^2 Write the quotient in standard form

OpenStudy (anonymous):

you will need so square the denomniator first and see what you get

OpenStudy (anonymous):

you should get \[(2+3i)^2=-5+12i\] so your next step is to divide \[\frac{5i}{-5+12i}\]and you do this by multiplying numerator and denominator by the conjugate of the denominator, in other words multiply by \[\frac{-5-12i}{-5-12I}\]

OpenStudy (anonymous):

the denominator will be \[5^2+12^2=169\] and the numerator will be \[5i(-5-12i)\] or \[-25i+60\] giving your final answer as \[\frac{60}{169}-\frac{25}{19}i\]

OpenStudy (anonymous):

typo on last line \[\frac{60}{169}-\frac{25}{169}i\]